OpenStudy (anonymous):

How do I put this in slope intercept form: -3x + 2y = 8 ?

OpenStudy (blues):

Solve for y by bringing all other terms to the other side of the equation. 2y = 8 + 3x y = 4 + 3/2 x Simple, really.

OpenStudy (anonymous):

-3x + 2y = 8 2y = 8 + 3x y = 1.5x + 4

OpenStudy (amistre64):

the question at this stage doesnt tend to be "how" to put it into slope int form; but rather, what is slope int form ....

OpenStudy (anonymous):

I've never referred to it as "slope intercept" form, but I'm assuming it's the method we learnt for constructing simple graphs back... whenever it was, where curves take the equation y = mx + c with m being the slope and c the intercept?

OpenStudy (amistre64):

yep, hence the name, slope intercept form

OpenStudy (blues):

I've never heard that called anything other than slope intercept form.

OpenStudy (anonymous):

The only curve sketching I've done in recent times involves calculus, so I don't recall what I've ever referred to it as - other than "y = mx + c." Not that I'm saying "slope intercept form" doesn't make sense or is a bad name.

OpenStudy (anonymous):

Any one of you care to explain how I can plot this?

OpenStudy (anonymous):

The intercept (in the context, the y-intercept) is where it crosses the y axis, and the slope is the value of m. So for every value of x, y is going to be m*x (starting at c). If that makes sense.

OpenStudy (blues):

The "c" is the value of the equation when x = 0, so you know that is the y intercept. The "x" is linear so you know that you're looking at a line, not a quadratic or some other polynomial. The coefficient in front of the x shows the "rise over the run." In totally non-mathematic terms.

OpenStudy (anonymous):

In this case the intercept (c) is 4, so plot a point on the y-axis at (0,4). Then - depending on your scale, but generally simple hand-drawn graphs increment by 1 - at x = 1, plot a point at 1.5x + 4 = 1.5*1 + 4 = 1.5 + 4 = 5.5 so (1,5.5). Then at x = 2, y = 1.5*2 + 4 = 3 + 4 = 7, so (2,7). And so on and so forth until you get as far as necessary.